calculate the pH of a saturated solution of magnesium hydroxide
1
Expert's answer
2020-01-15T06:56:03-0500

Magnesium hydroxide dissociates according to the following equation:

Mg(OH)2 ⇌ Mg2+ + 2OH-

Solubility product (Ksp) of magnesium hydroxide at standard conditions equals:

Ksp = 5.61 10¯12

As a result:

Ksp = [Mg2+] × [OH-]2

From here:

Ksp = (s) × (2s)2 = 4s3

s = (4Ksp)1/3 = 1.12 × 10-4 M

As a result, concentration of hydroxide ions equals:

[OH¯] = 2 × 1.12 × 10-4 M = 2.24 × 10-4 M

From here:

pOH = - log (2.24 × 10-4) = 3.650

pH = 14 - 3.650 = 10.350

Answer: 10.350